# Click Without Compromise: Online Advertising Measurement via Per User Differential Privacy

### v1

@@ 6.1. Baseline Methods
IPA (case2023interoperable, ): The Interoperable Private Attribution (IPA) protocol leverages a combination of multi-party computation and differential privacy to facilitate the processing of individuals’ data while ensuring that only aggregated measurements are disclosed. In this paper, our primary focus is on comparing how IPA achieves differential privacy, assuming the presence of a trusted third party that provides the joined table. IPA adds independent noise to the data vector ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, we use the group privacy property 2.8 to make sure it satisfies user-level neighboring datasets with a global contribution constraint of at most $GS$.
Id-BPC (add independent noise with bounded per-day contribution): We set a maximum number of contributions $r$ a user can make in 1 day, if any user contributes more than $r$ records (conversions) in 1 day, only the first $r$ records will be kept in the dataset, other records are cropped. Then we add independent noise to the data vector ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ (with proper scales). Id-BPC is proposed based on IPA to check the influence of bounded per day contributions.
@@
StreamUDP (dong2023continual, ): StreamUDP adds independent noise to a binary tree structure, we use the group privacy property 2.8 to make sure it satisfies user-level neighboring datasets with a global contribution constraint of at most $GS$.
Bin-BPC (binary tree method with bounded per-day contribution): Similar to Id-BPC, we set a maximum number of contributions $r$ a user can make in 1 day. Then we add independent noise to a binary tree structure (with proper scales). Bin-BPC is proposed based on StreamUDP to check the influence of bounded per day contributions.
@@
UMM (user-level matrix mechanism): It’s adapted from the event-level matrix mechanism (henzinger2022constant, ), we use the group privacy property 2.8 to make sure it satisfies user-level neighboring datasets with a global contribution constraint of at most $GS$.
6.2. Evaluation on Synthetic Datasets
In this section, we generate synthetic datasets to test the performance of our algorithm. We assume there are 1 million users, $|U|=10^{6}$. There are 1000 publishers and the privacy budget is $\rho=1$. For each user, we use 3 different distributions to generate the number of total conversions in the dataset (the distributions are similar to the settings in (dong2023continual, )). In this section, we assume last-touch attribution (LTA) is used. For example, if the distribution generates a number 5 for the user Alice, it means that Alice has 5 conversions in the dataset.
Since it’s improbable for a user to interact with all 1000 publishers, we opt to generate a random number $q$ within the range [1, 100] for each user, denoting the potential publishers they could interact with. Subsequently, we randomly select $q$ publishers from the pool of 1000 as the candidate set. For each conversion made by a user, we then randomly assign it to one of these $q$ publishers. Following this, we shuffle the dataset and evenly divide it into $n$ segments, where the i-th segment represents the conversions received on day i. Table 9 shows an example of the dataset. At each time (we assume each conversion happens in a non-overlapped timestamp), there is one conversion, that’s attributed to one of the publishers.
@@ 6.2. Evaluation on Synthetic Datasets
Table 9. Attribution weight for each publisher
For the algorithms StreamUDP and IPA, the global sensitivity of the dataset is denoted as $GS$. As for Bin-BPC, Id-BPC, and Ads-BPC, we set the threshold for daily contributions as $r$. Selecting an appropriate threshold value $r$ involves a trade-off between bias and variance. A small $r$ leads to the removal of numerous records, introducing substantial bias. Conversely, a large $r$ retains most records in the dataset but introduces significant noise into the results, resulting in high variance. In this paper, we manually determine the threshold $r$, while considering the adaptive selection of $r$ as a potential avenue for future research. The adaptive determination of $r$ is considered a parallel task to our current method, and any suitable technique (e.g., as discussed in (dong2023continual, )) can be applied to enhance our approach. We use $r=\lceil\frac{GS}{n}*2\rceil$ as the threshold, with the hope that most of the users have contributions less than $r$ on each day.
6.2.1. Experiment on Zipf Dataset
In the Zipf dataset, the number of conversions for each user is generated from the distribution: Zipf(3)+10. Here the global sensitivity is GS = 50 (A user has at most 50 conversions for each publisher). The total number of conversions in the dataset is $T=11,361,742$. The user contribution limit for each day is set as $r=\lceil\frac{50}{31}*2\rceil=4$. Table 10 displays the median relative error, 90% maximum relative error, and root mean squared error across different publishers for various algorithms. Ads-BPC stands out for achieving the lowest error among all algorithms. It demonstrates an improvement of approximately 35% over the second-best algorithm, Id-BPC.
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 1.61 1.20 2.66 0.83 1.70 0.55
90% Max Relative Error (%) 4.95 4.63 6.16 1.94 7.21 1.21
RMSE 126.08 94.72 174.71 54.53 127.27 37.62
Table 10. Error Comparison on Zipf Dataset
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 1.09 0.68 1.80 0.47 1.15 0.24
90% Max Relative Error (%) 3.35 2.61 4.18 1.10 4.87 0.54
RMSE 379.58 237.45 522.20 138.62 380.86 73.95
Table 11. Error Comparison on Normal Dataset
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 0.73 0.49 1.21 0.32 0.77 0.15
90% Max Relative Error (%) 2.24 1.97 2.82 0.84 3.27 0.36
RMSE 650.59 625.32 900.86 535.92 651.58 488.21
Table 12. Error Comparison on Uniform Dataset
6.2.2. Normal distribution
The number of conversions for each user is generated from the distribution: Normal(50, 30). Here the global sensitivity is GS = 150. The total number of conversions in the dataset is $T=50,193,976$. The user contribution limit for each day is $r=\lceil\frac{150}{31}*2\rceil=10$. In Table 11, Ads-BPC consistently outperforms other algorithms, showcasing the lowest median relative error, 90% maximum relative error, and root mean squared error. On average, Ads-BPC demonstrates a 48% improvement over Id-BPC, which ranks second in performance.
6.2.3. Uniform distribution
The number of conversions for each user is generated from the distribution: Uniform(1, 256). Here the global sensitivity is $GS=256$, The total number of conversions in the dataset is $T=51,180,247$. The user contribution limit for each day is $r=\lceil\frac{256}{31}*2\rceil=17$. Table 12 illustrates that Ads-BPC consistently outperforms other algorithms across various publishers, showcasing an average improvement of 50% over the second-best algorithm, Id-BPC.
6.3. Experiments on Criteo Dataset
The Criteo Sponsored Search Conversion Log Dataset (tallis2018reacting, ) contains logs obtained from Criteo Predictive Search (CPS). It offers an automated end-to-end solution using sophisticated machine learning techniques to enhance the Google Shopping experience through robust predictive optimization across all aspects of an advertiser’s campaign. Each row in the dataset represents a user action, such as a click, performed on a product-related advertisement. These product advertisements are displayed to users after they express intent through an online search engine. Each row in the dataset includes information about product characteristics (age, brand, gender, price), the timing of the click, user characteristics, and device information. Additionally, the logs contain information about whether these clicks eventually led to a conversion, which means the product was purchased within an $n$-day window, along with the time elapsed between the click and the conversion.
@@ 6.3. Experiments on Criteo Dataset
We use different attribution logic (LTA, FTA, UNI) in the experiment, and the results show that Ads-BPC performs the best regardless of the attribution logic. Table 13 presents the median relative error, 90% maximum relative error, and root mean squared error of different algorithms across various publishers when LTA is used. Notably, the Ads-BPC algorithm attains the lowest error compared to other algorithms, leading to an improvement of approximately 25% over Id-BPC. Table 14 shows the errors of different algorithms when FTA is used. Ads-BPC has the lowest error compared to other algorithms, with an improvement of approximately 30% over Id-BPC. Using UNI attribution logic, Table 15 presents the errors of different algorithms. Ads-BPC improves 33% over Id-BPC.
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 11.90 7.65 20.63 5.25 13.37 4.02
90% Max Relative Error (%) 38.15 30.14 48.35 12.69 57.37 9.34
RMSE 111.40 71.25 150.08 40.97 112.93 29.88
Table 13. Error Comparison on Criteo Dataset using LTA
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 7.78 5.02 12.06 3.37 8.05 2.38
90% Max Relative Error (%) 22.89 18.77 29.99 7.53 34.15 5.42
RMSE 112.51 70.42 154.09 40.80 111.85 30.37
Table 14. Error Comparison on Criteo Dataset using FTA
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 7.94 4.84 12.16 3.42 8.42 2.28
90% Max Relative Error (%) 23.78 18.84 30.05 8.18 36.10 5.41
RMS 111.09 71.31 151.98 41.16 112.39 30.13
Table 15. Error Comparison on Criteo Dataset using UNI
6.4. Experiments on Facebook Sales Conversion Dataset
@@
Facebook Sales Conversion Dataset (fbdata, ) is a dataset used for Facebook ad campaign analysis and sales prediction. Each row of the raw dataset contains information about age, gender, interest, and number of conversions from a unique user group. We treat each unique user group as a single user and distribute the conversions across $n$ days. Key dataset statistics include $\text{GS}=60$, $T=3264$, and $|U|=1143$. The conversions are distributed to $n=31$ days. The user contribution limit for each day is $r=\lceil\frac{60}{31}*2\rceil=4$. The privacy budget is $\rho=1$. Table 16 presents the median relative error, 90% maximum relative error, and root mean squared error of different algorithms across various publishers when LTA is used. Ads-BPC algorithm has an improvement of approximately 50% over the second best Id-BPC.
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 5.26 4.21 11.36 2.81 8.80 1.42
90% Max Relative Error (%) 18.75 15.46 27.48 6.68 31.99 3.27
RMSE 135.46 91.10 206.01 54.42 166.61 30.19
Table 16. Error Comparison on Facebook Sales Conversion Dataset
6.5. Experiments on Different Parameters
@@
In this section, we demonstrate how different parameters influence the errors. We use the Criteo dataset for illustration. The results for the Facebook Sales Conversion dataset are provided in the appendix.
6.5.1. Experiments on Different Ad Campaign Durations
Figure 1. Comparison of different number of days $n$ on Criteo Dataset.
Figure 2. Comparison of different privacy budgets $\rho$ on Criteo Dataset.
Figure 3. Comparison of different user contributions $r$ on Criteo Dataset.
In this section, we study the performance of different algorithms with respect to the number of days $n$, $n$ is chosen from [7, 15, 31, 63, 127, 255]. The dataset we use is the Criteo dataset with LTA as described in Section 6.3. The privacy budget is set as $\rho=1$. The user contribution limit for each day is $r=\lceil\frac{44}{n}*2\rceil$. Figure 1 illustrates the performance of various algorithms. Ads-BPC outperforms others consistently.
6.5.2. Experiment on Different Privacy Budgets
@@
In this section, we study the performance of different algorithms with respect to the privacy budget $\rho$, $\rho$ is chosen from [0.25, 0.5, 1, 2, 4, 8]. The dataset we use is the Criteo dataset with LTA as described in section 6.3. The number of days is $n=31$. The user contribution limit for each day is $r=\lceil\frac{44}{31}*2\rceil=3$. As depicted in Figure 2, the errors associated with various algorithms decrease as $\rho$ increases. This observed trend can be attributed to the fact that with a larger privacy budget, the amount of noise added to the dataset decreases, leading to lower error rates.
6.5.3. Experiments on Different User Contributions
@@
In this section, we study the performance of different algorithms with respect to the user contribution limits $r$, $r$ is chosen from [1, 2, 3, 4, 5, 6, 8, 10, 12]. The dataset we use is the Criteo dataset with LTA as described in Section 6.3. The number of days is $n=31$. The privacy budget is $\rho=1$. In Figure 3, as the user contribution limit increases, more noise is introduced into the data, resulting in higher variance. Ads-BPC consistently beats other baseline methods.
7. Conclusions
@@
Our work is the first to formulate the real-world problem of MTA advertising measurement systems with real-time reporting of streaming data in advertising campaigns. We introduced Ads-BPC, a novel differentially private solution in this context, which employs a matrix mechanism formulation and further minimizes the total noise power with an analytical solution, leading to non-i.i.d. noise distribution for streaming differentially private measurement results. Extensive experiments confirm that Ads-BPC significantly enhances measurement accuracy by 25% to 50% compared to existing streaming differential privacy mechanisms applied in advertising measurement. This improvement underscores the effectiveness of our approach in achieving superior accuracy while maintaining a formal privacy guarantee, thus advancing the state-of-the-art in privacy-preserving advertising measurement.
Acknowledgments
The authors used ChatGPT4o to revise the text in Section 4 and Section 6 to “Revise the following text, keep the latex syntax, fix typos and make it easy to understand".

### v2

@@ 6.2 Baseline Methods
IPA [10]: The Interoperable Private Attribution (IPA) protocol leverages a combination of multi-party computation and differential privacy to facilitate the processing of individuals’ data while ensuring that only aggregated measurements are disclosed. In this paper, our primary focus is on comparing how IPA achieves differential privacy, assuming the presence of a trusted third party that provides the joined table. IPA adds independent and identically distributed (IID) noise to the data vector ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, we use the group privacy property 2.9 to make sure it satisfies user-level neighboring datasets with a global contribution constraint of at most $GS$.
@@
BIN: BIN adds IID noise to a binary tree structure, we use the group privacy property 2.9 to make sure it satisfies user-level neighboring datasets with a global contribution constraint of at most $GS$.
Stream [19]: Similar to BIN, Stream adds IID noise to a binary tree structure. Rather than using a noise scale proportional to the global sensitivity $GS$, Stream employs the Sparse Vector Technique (SVT) to privately estimate the maximum number of contributions $\tau$ that a user can make in the dataset. Specifically, if a user exceeds $\tau$ contributions, Stream doubles the bound, updating $\tau\leftarrow 2\times\tau$.
@@
UMM (user-level matrix mechanism): It’s adapted from the event-level matrix mechanism [27], we use the group privacy property 2.9 to make sure it satisfies user-level neighboring datasets with a global contribution constraint of at most $GS$.
MMBPC (matrix mechanism with bounded per-day contribution): Similar to AdsBPC, we set a maximum contribution limit $r_{i}$ for each user on day $i$. If any user contributes more than $r_{i}$ conversions on a given day, only the first $r_{i}$ conversions are retained in the dataset. We first compute the event-level matrix mechanism and then scale the columns of the matrix based on the contribution limit $r_{i}$, ensuring that the sensitivity remains unchanged.
6.3 Datasets
We use three synthetic datasets and two real world datasets to test the performance of our algorithm. For synthetic datasets, we assume last-touch attribution (LTA) is used. For real datasets, last-touch attribution (LTA), first-touch attribution (FTA) and uniform atribution (UNI) are used to create different data vectors ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$. Table VIII shows an example of the dataset. At each time (we assume each conversion happens in a non-overlapped timestamp), there is one conversion, that’s attributed to one of the publishers.
@@ 6.3 Datasets
TABLE VIII: Attribution weight for each publisher
Dataset IPA BIN Stream UMM MMBPC AdsBPC Improvement
Zipf 42.54 43.79 129.69 37.89 45.63 21.09 44.33%
Normal 189.73 142.53 427.82 123.41 116.12 77.28 33.45%
Uniform 294.17 153.13 760.19 316.65 161.71 56.69 62.98%
Criteo (LTA) 42.86 36.70 68.97 38.29 30.84 18.23 40.88%
Criteo (FTA) 81.14 37.47 68.02 43.08 50.01 21.78 41.87%
Criteo (UNI) 58.27 35.27 63.98 34.92 41.85 17.02 51.26%
Facebook 46.93 34.30 74.25 37.52 20.99 9.78 53.41%
TABLE IX: The query is the prefix sum, the objective (error metric) is the weighted root mean squared error. The total privacy budget is $\rho=1$. The number of days is $n=31$.
Dataset IPA BIN Stream UMM MMBPC AdsBPC Improvement
Zipf 0.89 1.13 31.41 0.72 6.62 0.24 66.67%
Normal 29.72 19.19 266.40 16.65 111.83 1.35 91.89%
Uniform 32.66 171.18 1634.14 22.00 533.23 1.16 94.72%
Criteo (LTA) 0.90 2.86 15.01 1.44 23.95 0.13 85.56%
Criteo (FTA) 1.41 2.12 13.95 0.76 25.39 0.11 85.53%
Criteo (UNI) 1.00 2.32 7.02 0.69 7.52 0.12 82.61%
Facebook 0.89 4.40 10.86 1.12 0.26 0.06 76.92%
TABLE X: The query is the sliding window sum, the objective (error metric) is the max variance (divied by $10^{4}$). The total privacy budget is $\rho=1$. The number of days is $n=31$.
Figure 1: Comparison of different privacy budgets $\rho$ on Criteo Dataset. The number of days is $n=31$.
Figure 2: Comparison of different number of days $n$ on Criteo Dataset. The total privacy budget is $\rho=1$.
Synthetic Datasets For synthetic datasets, we assume there are 1 million users, $|U|=10^{6}$. There are 1000 publishers. For each user, we use 3 different distributions to generate the number of total conversions in the dataset (the distributions are similar to the settings in [19]). For example, if the distribution generates a number 5 for the user Alice, it means that Alice has 5 conversions in the dataset. We use the following three syntehtic datasets in the experiment. Zipf Dataset In the Zipf dataset, the number of conversions for each user is generated from the distribution: Zipf(3)+10. Here the global sensitivity is GS = 50 (A user has at most 50 conversions for each publisher). The total number of conversions in the dataset is $T=11,361,742$. Normal Dataset The number of conversions for each user is generated from the distribution: Normal(50, 30). Here the global sensitivity is GS = 150. The total number of conversions in the dataset is $T=50,193,976$. Uniform Dataset The number of conversions for each user is generated from the distribution: Uniform(1, 256). Here the global sensitivity is $GS=256$, The total number of conversions in the dataset is $T=51,180,247$.
Criteo Dataset The Criteo Sponsored Search Conversion Log Dataset [44] contains logs obtained from Criteo Predictive Search (CPS). Each row in the dataset represents a user action, such as a click, performed on a product-related advertisement, it includes information about product characteristics (age, brand, gender, price), the timing of the click, user characteristics, and device information.
@@
Facebook Sales Conversion Dataset Facebook Sales Conversion Dataset [1] is a dataset used for Facebook ad campaign analysis and sales prediction. Each row of the raw dataset contains information about age, gender, interest, and number of conversions from a unique user group. We treat each unique user group as a single user and distribute the conversions across $n$ days. Because Facebook is the only publisher, both FTA, LTA and UNI will give the same attribution weight. Key dataset statistics include $\text{GS}=60$, $T=3264$, and $|U|=1143$.
6.4 Experiments on Different Queries and Objective Functions
In this section, we compare AdsBPC with baseline methods across various datasets. For the Zipf, Normal, and Uniform datasets, we use LTA to generate the data vector ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$. For the Criteo and Facebook datasets, we generate three versions of the data vector ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ using different attribution methods: LTA, FTA, and UNI. We set the number of days to $n=31$ and the privacy budget to $\rho=1$.
In AdsBPC and MMBPC, we use the differentially private quantile algorithm for the first $l=7$ days, the quantile percentage is $p=99\%$, the bounding factor is $\Lambda=32$. In the SVT algorithm, the scale parameters are ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\uparrow}=1.3$ and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow}=0.8$, the thresholds are $T^{\uparrow}=T^{\downarrow}=50$, the maximum number of reports is $k=7$. For the allocation of privacy budget, we set $\rho_{1}=0.7\rho$, $\rho_{2}=0.15\rho/l$ and $\rho_{3}=0.15\rho$.
Prefix Sum Query with Weighted Root Mean Squared Error. Under Scenario 1 in Section 6.1, table IX presents the weighted root mean squared error (WRMSE) of various methods across different datasets. AdsBPC consistently outperforms other baseline methods. The final column displays the relative improvement over the second-best method; AdsBPC achieves 33% to 63% lower error.
Sliding Window Sum Query with Max Var Error. Under Scenario 2 in Section 6.1, table X presents the max variance error of various methods across different datasets. AdsBPC consistently outperforms other baseline methods. The final column displays the relative improvement over the second-best method; AdsBPC achieves 66% to 95% lower error.
We can see that AdsBPC outperforms the global sensitivity-based methods IPA, BIN, and UMM. The reason for this is that global sensitivity relies on the maximum number of contributions a user makes within the dataset, which is often significantly higher than the average contribution per user. As a result, these methods introduce an excessive amount of noise, leading to higher errors. In contrast, AdsBPC clips user contributions on a daily basis, effectively reducing the sensitivity and minimizing the added noise. Stream performs poorly as it is primarily designed for continuous releases exceeding $10^{5}$ events. In contrast, AdsBPC is better suited for advertising measurement, where the number of online releases is comparatively small. While MMBPC employs the same approach to calculate the per-day user contribution bound and utilizes a complex matrix factorization technique, its performance still falls short of AdsBPC. This highlights that AdsBPC is effective despite its simple design.
6.5 Experiments on Different Parameters
@@
In this section, we demonstrate how different parameters influence the errors. We use the Criteo dataset with LTA for illustration.
@@
Experiment on Different Privacy Budgets. In this section, we study the performance of different algorithms with respect to the privacy budget $\rho$, $\rho$ is chosen from [0.25, 0.5, 1, 2, 4, 8]. The dataset we use is the Criteo dataset with LTA as described in section 6.3. The number of days is $n=31$. As shown in Figure 1, the errors associated with various algorithms decrease as $\rho$ increases. This is because the amount of noise added to the dataset decreases with a larger privacy budget, leading to lower errors.
@@
Experiments on Different Ad Campaign Durations. In this section, we study the performance of different algorithms with respect to the number of days $n$, $n$ is chosen from [7, 15, 31, 63, 127]. The dataset we use is the Criteo dataset with LTA. The total privacy budget is set as $\rho=1$. Figure 2 illustrates the performance of various algorithms. AdsBPC outperforms others consistently.
7 Conclusions
@@
Our work is the first to formulate the real-world problem of MTA advertising measurement systems with real-time reporting of streaming data in advertising campaigns. We introduced AdsBPC, a novel differentially private solution for streaming advertising measurement results. Extensive experiments confirm that AdsBPC significantly enhances measurement accuracy by 33% to 95% compared to existing streaming differential privacy mechanisms applied in advertising measurement. This improvement underscores the effectiveness of our approach in achieving superior accuracy while maintaining a formal privacy guarantee, thus advancing the state-of-the-art in privacy-preserving advertising measurement.